OpenStudy (ledah):

solving systems of equations 2x+y=-6 3x+y=-10

OpenStudy (diyadiya):

Subtract both equations

OpenStudy (diyadiya):

2x+y=-6 3x+y=-10 ---------- -x +0 =4 so -x =4 x=-4 Now plug in the value of x in any of the equations and you'll get the value of y

OpenStudy (ledah):

okay I'll do that now

OpenStudy (ledah):

is it 2?

OpenStudy (diyadiya):

2(-4) + y =-6 -8 + y = -6 y=-6+8 y=2 Yes! YOu're Right :D

OpenStudy (ledah):

thanks alot :D

OpenStudy (diyadiya):

Anytime :)